OpenStudy (anonymous):

Which of the following, if any, is the multiplicative inverse of 9/11? A) A) 1 2/9 B) B) 9/11 C) C) The multiplicative inverse does not exist. D) D) None of the other choices.

OpenStudy (anonymous):

On your calculator, do 9/11 Take that decimal and raise it to -1

OpenStudy (anonymous):

Then convert that number back into fraction

OpenStudy (anonymous):

so then -.0909 back into a fraction

OpenStudy (anonymous):

9/11 should equal .8181 raising that to -1 should equal 1.222222

OpenStudy (anonymous):

Which is 61/50

OpenStudy (anonymous):

It doesn't exist

Directrix (directrix):

A number and its multiplicative inverse have a product of 1. Multiplicative inverse of a number is another name for reciprocal. The reciprocal of 9/11 DOES exist. What times 9/11 = 1?

Directrix (directrix):

@schamp2010 What is the reciprocal of 9/11?

OpenStudy (anonymous):

-11/9

OpenStudy (anonymous):

Sorry about that Schamp. I was thinking of something else.

Directrix (directrix):

You have additive inverse and multiplicative inverse mixed up.

Directrix (directrix):

A number and its additive inverse add to zero. Example: 5 + (-5) = 0 so -5 is the additive inverse of 5.

Directrix (directrix):

But, we are not doing that. We are doing multiplicative inverses.

OpenStudy (anonymous):

Answer should be a number that will cancel the numerator and denominator \[\frac{ 9 }{ 11 }*\frac{ x }{ y}\]

OpenStudy (anonymous):

So the answer isn't there

OpenStudy (anonymous):

It is there. What number divided by ( will give you 1?

Directrix (directrix):

Multiplicative inverses multiply to ONE. @schamp2010 I SEE the answer. If you tell me what the reciprocal of 9/11 is, I will show you that the answer is listed among the options.

OpenStudy (anonymous):

\[\frac{ x }{ 11 }=1 and \frac{ y }{ 9 }\]

OpenStudy (anonymous):

11/9 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Postive 11/9

OpenStudy (anonymous):

So then the answer is a?

OpenStudy (anonymous):

You listed A) is 12/9. Did you copy thast correctly?

OpenStudy (anonymous):

1 2/9

Directrix (directrix):

A) 1 2/9 is 1 and 2/9. Change 1 to 9/9. Add 9/9 + 2/9 to get 11/9 which is the multiplicative inverse of 9/11.